# Approaches to critical point theory via sequential and parametrized topological complexity

### v2

@@ 6. Vertically proportional vector fields and parametrized subspace complexities
Note that $E$ and $X$ are path-connected ANRs by [Pal66a, Cor. to Theorem 5]. Since $p^{r}\colon E^{r}\to X^{r}$ is submersive as well, it follows that
$E_{X}^{r}=(p^{r})^{-1}(\Delta_{r,X})$
is a smooth Banach manifold whose tangent space at $u=(u_{1},\dots,u_{r})\in E^{r}_{X}$ is given by
@@
$\mathrm{Ver}_{u}(q_{r})$ =$=\{v\in T_{u}E^{r}_{X}\mid(Dq_{r})_{u}(v)=0\}$
=$=\{(v_{1},\dots,v_{r})\in T_{u}E^{r}\mid 0=Dp_{u_{1}}(v_{1})=Dp_{u_{2}}(v_{2})=\dots=Dp_{u_{r}}(v_{r})\}$
$=\mathrm{Ver}_{u}(p^{r}).$
@@
    We define
    $\mathrm{pr}_{\mathrm{ver}}\colon TM\to TM$
    as the bundle map for which $\mathrm{pr}_{\mathrm{ver}}|_{T_{m}M}$ is the orthogonal projection onto $\mathrm{Ver}_{m}(q)$ with respect to $g$ for each $m\in M$.
@@
    Given $F\in C^{1}(M)$, we further put
    $\nabla^{\mathrm{ver}}F:=(\nabla^{g}F)^{\mathrm{ver}},$
    where $\nabla^{g}F\colon M\to TM$ denotes the gradient of $F$ with respect to $g$.
@@
Let $M$, $N$, $g$ and $q$ be given as in Definition 6.2 and let $Y\colon M\to TM$ be a vector field. Apparently, $\mathrm{sing}(Y)\subset\mathrm{sing}(Y^{\mathrm{ver}})$. Conversely, if $Y$ is vertically proportional, then every singular point of $Y^{\mathrm{ver}}$ is a singular point of $Y$ as well. Thus, if $Y$ is vertically proportional, then it necessarily holds that
@@
As we will see in the next results, if a function $F\in C^{1}(E^{r}_{X})$ has a vertically proportional gradient with respect to a complete Riemannian metric, then this will allow us to apply Theorem 6.1 to estimate the number of critical points of $F$ by sequential parametrized subspace complexities. However, as condition (6.3) already indicates, the vertical proportionality of its gradient is a strong restriction on the function itself.
@@
    Let $\phi^{\mathrm{ver}}_{1}\colon M\to M$ be the time-1 map of the flow of $-Y^{\mathrm{ver}}$. Then $(\phi^{\mathrm{ver}}_{1},F)$ satisfies condition (D).
@@
[…]
@@
      Evidently, $\|Y^{\mathrm{ver}}(m)\|\leq\|Y(m)\|$ for each $m\in M$, so the upper bound on $\|Y^{\mathrm{ver}}(m)\|$ follows immediately from the corresponding upper bound on $\|Y\|$, which exists since $Y$ is a pseudo-gradient.
@@
      One easily sees that
      $Y^{\mathrm{ver}}(m)=h(m)\cdot\nabla^{\mathrm{ver}}F(m)\qquad\forall m\in M.$
      Thus, we compute for each $m\in M$ that
@@
      $=h(m)\cdot g_{M}(\nabla^{\mathrm{ver}}F(m),\nabla^{\mathrm{ver}}F(m))=h(m)\cdot\|\nabla^{\mathrm{ver}}F(m)\|^{2}$
      $\geq h(m)\cdot C^{-2}\|\nabla^{g}F(m)\|^{2}.$
      If $\|\nabla^{g}F(m)\|\leq 1$, then it follows immediately that
      $DF_{m}(Y^{\mathrm{ver}}(m))\geq C^{-2}\min\{\|\nabla^{g}F(m)\|,\|\nabla^{g}F(m)\|^{2}\}.$
      For all $m\in M$ with $\|\nabla^{g}F(m)\|>1$ it holds that $h(m)\geq\frac{1}{\|\nabla^{g}F(m)\|}$ and thus
      $DF_{m}(Y^{\mathrm{ver}}(m))\geq C^{-2}\|\nabla^{g}F(m)\|=C^{-2}\min\{\|\nabla^{g}F(m)\|,\|\nabla^{g}F(m)\|^{2}\}.$
      This shows property (iii) of a pseudo-gradient if we put $C_{2}:=C^{-2}$ in the situation of Definition 1.5.
@@
Throughout the rest of this section, we choose and fix a complete Riemannian metric $g$ on $E$ and let $\bar{g}$ be the Riemannian metric on $E^{r}_{X}$ that is obtained by restricting the product metric $g^{r}$ on $E^{r}$.
@@
This follows from Theorem 6.5 in the same way as Theorem 3.3 is derived from Corollary 2.11, since the sequential motion planner by constant paths on $\Delta_{r,E}$ is indeed a continuous sequential parametrized motion planner which extends to an open neighborhood of $\Delta_{r,E}$ by Lemma 5.8. ∎
@@
[…]
@@
    with respect to the Euclidean inner product. With $\mathrm{pr}_{2}\colon\mathbb{R}^{N_{1}}\times\mathbb{R}^{2N_{2}}\to\mathbb{R}^{2N_{2}}$ denoting the projection, we observe that
    $F=\varphi\circ\mathrm{pr}_{2}\circ i_{0}$
    and thus
@@
    Let $(e_{1},e_{2})\in E\times_{B}E$ and let $u\in\mathrm{Hor}_{(e_{1},e_{2})}(q)$. We write $(Di_{0})_{(e_{1},e_{2})}(u)=(u_{0},u_{1},u_{2})\in\mathbb{R}^{n_{1}+2n_{2}}$. Set $(x_{1},x_{2})=(i_{E}(e_{1}),i_{E}(e_{2}))\in\mathbb{R}^{2n_{2}}$. Then equation (6.5) yields
    $DF_{(e_{1},e_{2})}(u)=D\varphi_{(x_{1},x_{2})}(u_{1},u_{2})=2\langle x_{1}-x_{2},u_{1}-u_{2}\rangle,$
@@
Note that condition (6.4) in part b) of Proposition 6.9 is satisfied if for all $(e_{1},e_{2})\in E\times_{B}E$ and $u\in\mathrm{Hor}_{(e_{1},e_{2})}(q)$ we have that $(Di_{0})_{(e_{1},e_{2})}(u)\in\mathbb{R}^{n_{1}}\times\Delta_{\mathbb{R}^{n_{2}}}$, i.e. it is of the form $(Di_{0})_{(e_{1},e_{2})}(u)=(u_{0},u_{1},u_{1})$ for some $u_{0}\in\mathbb{R}^{n_{1}}$ and some $u_{1}\in\mathbb{R}^{n_{2}}$.
@@ 7. Sequential parametrized TC of restrictions of fibrations to sublevel sets
Here, we want to derive upper bounds on $\operatorname{\mathsf{TC}}[p|_{f^{\lambda}}\colon f^{\lambda}\to X]$ under the fibration assumption, i.e. upper bounds on sequential parametrized TCs instead of sequential parametrized subspace complexities. Such a result can be seen as a “parametrized analogue” of Corollary 4.3 on sublevel sets.
@@
The inclusion“$\supset$” is evident, so it only remains to show that every critical point of $F$ arises in this way. Let $u=(u_{1},\dots,u_{r})\in\operatorname{Crit}F$, such that
@@
$\nabla^{\mathrm{ver}}F(u)=\left(\nabla^{\mathrm{ver}}f(u_{1}),,\dots,\nabla^{\mathrm{ver}}f(u_{r})\right)\quad\forall u=(u_{1},\dots,u_{r})\in E^{r}_{X},$
@@
$\left(\mathrm{Ver}_{u}(q_{r})\right)^{\perp}=\bigcap_{i=1}^{r}E_{i}^{\perp}.$
Apparently, $E_{i}^{\perp}=T_{u}E^{r}_{X}\cap(\prod_{j=1}^{i-1}T_{u_{j}}E\times(\mathrm{Ver}_{u_{i}}(p))^{\perp}\times\prod_{j=i+1}^{r}T_{u_{j}}E)$ for each $i\in\{1,2,\dots,r\}$, such that
$\left(\mathrm{Ver}_{u}(q_{r})\right)^{\perp}=T_{u}E^{r}_{X}\cap\prod_{i=1}^{r}(\mathrm{Ver}_{u_{i}}(p))^{\perp}.$
@@
By definition, $F=f^{r}\circ i$, where $i\colon E^{r}_{X}\hookrightarrow E^{r}$ denotes the embedding, and one checks without difficulties that
@@
$\|\nabla^{\bar{g}}F(u)\|_{\bar{g}}=\|(\nabla^{g}f(u_{1}),\dots,\nabla^{g}f(u_{r}))\|_{g^{r}}=\sqrt{\sum_{i=1}^{r}\|\nabla^{g}f(u_{i})\|_{g}^{2}},$
where the norms are induced by the metrics in their subscripts. Since $\nabla f$ is vertically proportional, there exists $C>0$ with $\|\nabla f(e)\|_{g}\leq C\cdot\|\nabla^{\mathrm{ver}}f(e)\|_{g}$ for all $e\in E$. Thus,
$\|\nabla F(u_{1},\dots,u_{r})\|_{\bar{g}}$ $\leq\sqrt{C^{2}\sum_{i=1}^{r}\|\nabla^{\mathrm{ver}}f(u_{i})\|_{g}^{2}}=C\cdot\|(\nabla^{\mathrm{ver}}f(u_{1}),\dots,\nabla^{\mathrm{ver}}f(u_{r}))\|_{g^{r}}$
$=C\cdot\|\nabla^{\mathrm{ver}}F(u_{1},\dots,u_{r})\|_{\bar{g}},$
@@
$\operatorname{Crit}_{\mu}F$ $=F_{\mu}\cap\left(\operatorname{Crit}f\times_{X}\operatorname{Crit}f\times_{X}\dots\times_{X}\operatorname{Crit}f\right)$
$=\bigcup_{\mu_{1}+\mu_{2}+\dots+\mu_{r}=\mu}\operatorname{Crit}_{\mu_{1}}f\times_{X}\operatorname{Crit}_{\mu_{2}}f\times_{X}\dots\times_{X}\operatorname{Crit}_{\mu_{r}}f.$
@@
We recall that in Theorem 4.2 we have established an upper bound on sequential TCs of sublevel sets by sequential subspace complexities that are not taken with respect to the whole domain of the function, but only with respect to the sublevel set itself. We want to establish an analogue of this result for sequential parametrized TCs and restrictions of fibrations of the form $p_{\lambda}:=p|_{f^{\lambda}}\colon f^{\lambda}\to X$. However, $p_{\lambda}$ will in general not be a fibration and will only have this property in very particular settings.
Remark 7.7.
The parametrized TC of maps which are not fibrations was studied in a general way by García Calcines in [GC23]. However, it remains an open question how much of the topological Lusternik-Schnirelmann theory that we used to establish our main estimates can be transferred to the more general setting.
We refer the reader to the next section for examples of situations in which this assumption is indeed satisfied.
@@
Let $Y_{f}\colon E\to TE$ and $Y_{F}\colon E^{r}_{X}\to TE^{r}_{X}$ be the pseudo-gradients of $f$ and $F$, respectively, that are constructed as in Proposition 6.4.a). Let $\varphi^{\mathrm{ver}}\colon E\times\mathbb{R}\to E$ and $\Phi^{\mathrm{ver}}\colon E^{r}_{X}\times\mathbb{R}\to E^{r}_{X}$ denote the flows of $Y_{f}^{\mathrm{ver}}$ and $Y_{F}^{\mathrm{ver}}$, respectively. One derives from Lemma 7.5 that
$Y_{F}^{\mathrm{ver}}(u_{1},\dots,u_{r})=(Y_{f}^{\mathrm{ver}}(u_{1}),\dots,Y_{f}^{\mathrm{ver}}(u_{r}))$
and thus
@@
Since $f(\operatorname{Crit}f)$ is discrete, we can assume w.l.o.g. that $(\lambda,\lambda+\delta]$ does not contain a critical value of $f$. Put $U:=(f^{<\lambda+\delta})^{r}\cap E^{r}_{X}$. One derives from (7.2) that $\Phi^{\mathrm{ver}}(u,t)\in U$ for all $u\in U$ and $t\in[0,+\infty)$, such that $\Phi^{\mathrm{ver}}$ restricts to a flow
$\Psi:=\Phi^{\mathrm{ver}}|_{U\times[0,+\infty)}\colon U\times[0,+\infty)\to U.$
In other words, $Y_{F}^{\mathrm{ver}}|_{U}$ is forward complete. By Proposition 6.4.b), $(\varphi^{\mathrm{ver}}_{1},f)$ satisfies condition (D). Since $\partial U\subset(f_{\lambda+\delta})^{r}$ and since $\lambda+\delta$ is a regular value of $f$, it holds that $\partial U\cap\mathrm{Fix\;}\Phi=\emptyset$. Thus, by Lemma 4.1.b) and (7.2), $(\Psi_{1},F|_{U})$ satisfies condition (D).
@@
$\operatorname{\mathsf{TC}}_{r}[p_{\lambda}\colon f^{\lambda}\to B]$ $=\operatorname{\mathsf{TC}}_{r,p_{\lambda}}((f^{\lambda})^{r}_{X})=\operatorname{\mathsf{TC}}_{r,p_{\lambda}}\left((F|_{U})^{r\lambda}\right)$
$\leq\sum_{\mu\in(-\infty,r\lambda]}\operatorname{\mathsf{TC}}_{r,p_{\lambda}}(F_{\mu}\cap\operatorname{Crit}F).$

### v4

@@ 6. Vertically proportional vector fields and parametrized subspace complexities
Note that $E$ and $X$ are path-connected ANRs by [Pal66a, Cor. to Theorem 5]. Since $p^{r}\colon E^{r}\to X^{r}$ is submersive as well, it follows that $E_{X}^{r}=(p^{r})^{-1}(\Delta_{r,X})$ is a smooth Banach manifold whose tangent space at $u=(u_{1},\dots,u_{r})\in E^{r}_{X}$ is given by
@@
$\mathrm{Ver}_{u}(q_{r})$ $=\{v\in T_{u}E^{r}_{X}\mid(Dq_{r})_{u}(v)=0\}$
$=\{(v_{1},\dots,v_{r})\in T_{u}E^{r}\mid 0=Dp_{u_{1}}(v_{1})=Dp_{u_{2}}(v_{2})=\dots=Dp_{u_{r}}(v_{r})\}=\mathrm{Ver}_{u}(p^{r}).$
@@
    Let $\mathrm{pr}_{\mathrm{ver}}\colon TM\to TM$ be the bundle map for which $\mathrm{pr}_{\mathrm{ver}}|_{T_{m}M}$ is the orthogonal projection onto $\mathrm{Ver}_{m}(q)$ with respect to $g$ for each $m\in M$.
@@
    Given $F\in C^{1}(M)$, we further put $\nabla^{\mathrm{ver}}F:=(\nabla^{g}F)^{\mathrm{ver}}$, where $\nabla^{g}F\colon M\to TM$ denotes the gradient of $F$ with respect to $g$.
@@
Let $M$, NN, $g$ and $q$ be given as in Definition 6.2 and let $Y\colon M\to TM$ be a vector field. Apparently, $\mathrm{sing}(Y)\subset\mathrm{sing}(Y^{\mathrm{ver}})$. Conversely, if $Y$ is vertically proportional, then every singular point of $Y^{\mathrm{ver}}$ is a singular point of $Y$ as well. Thus, if $Y$ is vertically proportional, then it necessarily holds that
@@
As we will see in the next results, the existence of a vertical proportional gradient of a function $F\in C^{1}(E^{r}_{X})$ will allow us to apply Theorem 6.1 to estimate the number of critical points of $F$ by sequential parametrized subspace complexities. However, as condition (6.3) already indicates, the vertical proportionality of its gradient is a strong restriction on the function itself.
@@
    Let $\phi^{\mathrm{ver}}_{1}\colon M\to M$ be the time-1 map of the flow of $Y^{\mathrm{ver}}$. Then $(\phi^{\mathrm{ver}}_{1},F)$ satisfies condition (D).
@@
As discussed in Example 1.7, $Y$ is a pseudo-gradient for $F$. For each $m\in M$ we let $\|\cdot\|$ denote the norm on $T_{m}M$ induced by $g$. By assumption, there exists $C\in(0,+\infty)$, such that $\|\nabla^{g}F(m)\|\leq C\cdot\|\nabla^{\mathrm{ver}}F(m)\|$ for all $m\in M$. By definition of $Y$, this immediately yields that
@@
      Evidently, $\|Y^{\mathrm{ver}}(m)\|\leq\|Y(m)\|$ for each $m\in M$, so the upper bound of $\|Y^{\mathrm{ver}}(m)\|$ follows immediately from the corresponding upper bound of $\|Y\|$, which exists since $Y$ is a pseudo-gradient.
@@
      One easily sees that $Y^{\mathrm{ver}}(m)=h(m)\cdot\nabla^{\mathrm{ver}}F(m)$ for each $m\in M$. Thus,
@@
      $=h(m)\cdot\|\nabla^{\mathrm{ver}}F(m)\|^{2}\geq h(m)\cdot C^{-2}\|\nabla^{g}F(m)\|^{2}$
      for each $m\in M$. One further computes that
      $DF_{m}(Y^{\mathrm{ver}}(m))\geq C^{-2}\|\nabla^{g}F(m)\|=C^{-2}\min\{\|\nabla^{g}F(m)\|,\|\nabla^{g}F(m)\|^{2}\}\qquad\forall m\in M,$
      which shows property (iii) of a pseudo-gradient if we put $C_{2}:=C^{-2}$ in Definition 1.5.
@@
Throughout the rest of this section, we fix a complete Riemannian metric $g$ on $E$ and let $\bar{g}$ be the Riemannian metric on $E^{r}_{X}$ that is obtained by restricting the product metric $g^{r}$ on $E^{r}$.
@@
This follows from Theorem 6.5 in the same way as Theorem 3.3 is derived from Corollary 2.9. The sequential motion planner by constant paths on $\Delta_{r,E}$ is indeed a sequential parametrized motion planner which extends to an open neighborhood of $\Delta_{r,E}$ by Lemma 5.9. ∎
@@
where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{N_{1}}$.
@@
    with respect to the Euclidean inner product. With $\mathrm{pr}_{2}\colon\mathbb{R}^{N_{1}}\times\mathbb{R}^{2N_{2}}\to\mathbb{R}^{2N_{2}}$ denoting the projection, we observe that $F=\varphi\circ\mathrm{pr}_{2}\circ i_{0}$ and thus
@@
    Let $(e_{1},e_{2})\in E\times_{B}E$ and $u\in\mathrm{Hor}_{(e_{1},e_{2})}(q)$. We write $(Di_{0})_{(e_{1},e_{2})}(u)=(u_{0},u_{1},u_{2})\in\mathbb{R}^{n_{1}+2n_{2}}$. Then equation (6.5) yields
    $DF_{(e_{1},e_{2})}(u)=D\varphi_{(x_{1},x_{2})}(u_{1},u_{2})=2\langle i_{E}(e_{1})-i_{E}(e_{2}),u_{1}-u_{2}\rangle,$
@@
Condition (6.4) in Proposition 6.9.b) is satisfied if $(Di_{0})_{(e_{1},e_{2})}(u)\in\mathbb{R}^{n_{1}}\times\Delta_{\mathbb{R}^{n_{2}}}$ for all $(e_{1},e_{2})\in E\times_{B}E$ and $u\in\mathrm{Hor}_{(e_{1},e_{2})}(q)$, i.e. if $(Di_{0})_{(e_{1},e_{2})}(u)=(u_{0},u_{1},u_{1})$ for some $u_{0}\in\mathbb{R}^{n_{1}}$ and some $u_{1}\in\mathbb{R}^{n_{2}}$.
Example 6.11.
Let $k,n\in\mathbb{N}$ with $2\leq k\leq n-1$. Consider the Stiefel manifold of orthonormal $k$-frames in $\mathbb{R}^{n}$
$V_{k}(\mathbb{R}^{n})=\{X\in M_{n,k}(\mathbb{R})\,|\,X^{T}X=1_{k}\}$
where $M_{n,k}(\mathbb{R})$ denotes the set of real $n\times k$-matrices and $1_{k}\in M_{k,k}(\mathbb{R})$ is the identity matrix. For $X\in M_{n,k}(\mathbb{R})$ we write $X=(x_{1},\ldots,x_{k})$ where $x_{1},\ldots,x_{k}\in\mathbb{R}^{n}$ are the columns of $X$. We see that $V_{k}(\mathbb{R}^{n})$ is an embedded submanifold of $M_{n,k}(\mathbb{R})\cong\mathbb{R}^{nk}$ and we endow $V_{k}(\mathbb{R}^{n})$ with the Riemannian metric induced by the Euclidean metric on $\mathbb{R}^{nk}$. Then
$p\colon V_{k}(\mathbb{R}^{n})\to\mathbb{S}^{n-1},\qquad p(X)=x_{1}\quad\text{for}\ \ X=(x_{1},\ldots,x_{k})\in V_{k}(\mathbb{R}^{n}),$
is a fiber bundle with typical fiber $V_{k-1}(\mathbb{R}^{n})$. Proposition 6.9 yields a parametrized $2$-navigation function $F\colon V_{k}(\mathbb{R}^{n})\times_{\mathbb{S}^{n-1}}V_{k}(\mathbb{R}^{n})\to\mathbb{R}$ given by
$F(X,Y)=\|X-Y\|^{2}\quad\text{for }\quad(X,Y)\in V_{k}(\mathbb{R}^{n})\times_{\mathbb{S}^{n-1}}V_{k}(\mathbb{R}^{n}).$
Here $\|\cdot\|$ denotes the Euclidean norm on $M_{n,k}(\mathbb{R})\cong\mathbb{R}^{nk}$. For $X=(x_{1},\ldots,x_{k}),Y=(y_{1},\ldots,y_{k})\in V_{k}(\mathbb{R}^{n})$ with $p(X)=p(Y)$, it holds that $x_{1}=y_{1}$. Thus, we see that
$F(X,Y)=\sum_{j=2}^{k}\|x_{j}-y_{j}\|^{2}\qquad\forall(X,Y)\in V_{k}(\mathbb{R}^{n})\times_{\mathbb{S}^{n-1}}V_{k}(\mathbb{R}^{n}).$
A computation of the horizontal space of the bundle $V_{k}(\mathbb{R}^{n})\times_{\mathbb{S}^{n-1}}V_{k}(\mathbb{R}^{n})\to\mathbb{S}^{n-1}$ shows that we are in the situation of Remark 6.10 and the condition of Proposition 6.9.b) is satisfied in this example. The gradient vector field $\nabla^{g_{0}}F$ is thus vertically proportional.
@@ 7. Sequential parametrized TC of restrictions of fibrations to sublevel sets
Here, we want to derive upper bounds of $\operatorname{\mathsf{TC}}[p|_{f^{\lambda}}\colon f^{\lambda}\to X]$ under the fibration assumption, i.e. upper bounds of sequential parametrized TCs instead of sequential parametrized subspace complexities. Such a result can be seen as a “parametrized analogue” of Corollary 4.3 on sublevel sets.
@@
The inclusion $\supset$” is evident, so it only remains to show that every critical point of $F$ arises in this way. Let $u=(u_{1},\dots,u_{r})\in\operatorname{Crit}F$, such that
@@
$\nabla^{\mathrm{ver}}F(u)=\left(\nabla^{\mathrm{ver}}f(u_{1}),\dots,\nabla^{\mathrm{ver}}f(u_{r})\right)\quad\forall u=(u_{1},\dots,u_{r})\in E^{r}_{X},$
@@
$\left(\mathrm{Ver}_{u}(q_{r})\right)^{\perp}$ $=\bigcap_{i=1}^{r}E_{i}^{\perp}=\bigcap_{i=1}^{r}\Big(T_{u}E^{r}_{X}\cap\Big(\prod_{j=1}^{i-1}T_{u_{j}}E\times(\mathrm{Ver}_{u_{i}}(p))^{\perp}\times\prod_{j=i+1}^{r}T_{u_{j}}E\Big)\Big)$
$=T_{u}E^{r}_{X}\cap\prod_{i=1}^{r}(\mathrm{Ver}_{u_{i}}(p))^{\perp}.$
@@
Since $\nabla^{g}f$ is vertically proportional, there exists $C>0$ with $\|\nabla^{g}f(e)\|_{g}\leq C\cdot\|\nabla^{\mathrm{ver}}f(e)\|_{g}$ for all $e\in E$. By definition, $F=f^{r}\circ i$, where $i\colon E^{r}_{X}\hookrightarrow E^{r}$ denotes the embedding, and one checks without difficulties that
@@
$\|\nabla^{\bar{g}}F(u)\|_{\bar{g}}$ $=\|(\nabla^{g}f(u_{1}),\dots,\nabla^{g}f(u_{r}))\|_{g^{r}}=\sqrt{\sum_{i=1}^{r}\|\nabla^{g}f(u_{i})\|_{g}^{2}}$
$\leq\sqrt{C^{2}\sum_{i=1}^{r}\|\nabla^{\mathrm{ver}}f(u_{i})\|_{g}^{2}}=C\cdot\|\nabla^{\mathrm{ver}}F(u_{1},\dots,u_{r})\|_{\bar{g}},$
@@
$\operatorname{Crit}_{\mu}F$ $=\bigcup_{\mu_{1}+\mu_{2}+\dots+\mu_{r}=\mu}\operatorname{Crit}_{\mu_{1}}f\times_{X}\operatorname{Crit}_{\mu_{2}}f\times_{X}\dots\times_{X}\operatorname{Crit}_{\mu_{r}}f.$
@@
We recall that in Theorem 4.2 we have established an upper bound of sequential TCs of sublevel sets by sequential subspace complexities that are not taken with respect to the whole domain of the function, but only with respect to the sublevel set itself. We want to establish an analogue of this result for sequential parametrized TCs and restrictions of fibrations of the form $p_{\lambda}:=p|_{f^{\lambda}}\colon f^{\lambda}\to X$. However, $p_{\lambda}$ will in general not be a fibration and will only have this property in very particular settings. We refer the reader to the next section for examples of situations in which this assumption is indeed satisfied.
@@
Let $Y_{f}\colon E\to TE$ and $Y_{F}\colon E^{r}_{X}\to TE^{r}_{X}$ be the pseudo-gradients of $f$ and $F$, respectively, that are constructed as in Proposition 6.4.a). Let $\varphi^{\mathrm{ver}}\colon E\times\mathbb{R}\to E$ and $\Phi^{\mathrm{ver}}\colon E^{r}_{X}\times\mathbb{R}\to E^{r}_{X}$ denote the flows of $Y_{f}^{\mathrm{ver}}$ and $Y_{F}^{\mathrm{ver}}$, respectively. One derives from Lemmas 7.4 and 7.5 that
@@
Since $f(\operatorname{Crit}f)$ is discrete, we can assume w.l.o.g. that $(\lambda,\lambda+\delta]$ does not contain a critical value of $f$. Put $U:=(f^{<\lambda+\delta})^{r}\cap E^{r}_{X}$. One derives from (7.2) that $\Phi^{\mathrm{ver}}(u,t)\in U$ for all $u\in U$ and $t\in[0,+\infty)$, such that $\Phi^{\mathrm{ver}}$ restricts to a flow $\Psi:=\Phi^{\mathrm{ver}}|_{U\times[0,+\infty)}\colon U\times[0,+\infty)\to U$. In other words, $Y_{F}^{\mathrm{ver}}|_{U}$ is forward complete. By Proposition 6.4.b), $(\varphi^{\mathrm{ver}}_{1},f)$ satisfies condition (D). Since $\partial U\subset(f_{\lambda+\delta})^{r}$ and since $\lambda+\delta$ is a regular value of $f$, it holds that $\partial U\cap\mathrm{Fix\;}\Phi=\emptyset$. Thus, by Lemma 4.1.b) and (7.2), $(\Psi_{1},F|_{U})$ satisfies condition (D).
@@
$\operatorname{\mathsf{TC}}_{r}[p_{\lambda}\colon f^{\lambda}\to B]$ $=\operatorname{\mathsf{TC}}_{r,p_{\lambda}}((f^{\lambda})^{r}_{X})=\operatorname{\mathsf{TC}}_{r,p_{\lambda}}\left((F|_{U})^{r\lambda}\right)\leq\sum_{\mu\in(-\infty,r\lambda]}\operatorname{\mathsf{TC}}_{r,p_{\lambda}}(F_{\mu}\cap\operatorname{Crit}F).$